# Continual Release Moment Estimation with Differential Privacy

### v1

@@ Abstract
We propose Joint Moment Estimation (JME), a method for continually and privately estimating both the first and second moments of data with reduced noise compared to naive approaches. JME uses the matrix mechanism and joint sensitivity analysis to allow the second-moment estimation with no additional privacy cost, thereby improving accuracy while maintaining privacy. We demonstrate JME’s effectiveness in two applications: estimating the running mean and covariance matrix for Gaussian density estimation and model training with DP-Adam on CIFAR-10.
Machine Learning, ICML
@@ 1 Introduction
Estimating the first and second moments of data is a fundamental step in many machine learning algorithms, ranging from foundational methods, such as linear regression, principal component analysis, or Gaussian model fitting, to state-of-the-art neural network components, such as BatchNorm (Ioffe & Szegedy, 2015), and optimizers, such as Adam (Diederik & Ba, 2015). Often, these estimates must be computed and updated continuously, called the continual release setting (Dwork et al., 2010). For example, this is the case when the data arrives sequentially and intermediate results are needed without delay, such as for sequential optimization algorithms, such as Adam, or for real-time systems in healthcare, finance, or online recommendation systems. In such scenarios, ensuring the privacy of sensitive data, such as user information or medical records, is essential.
@@
Differential Privacy (DP) is a widely established formal notion of privacy that ensures that the inclusion or exclusion of any single individual’s data has a limited impact on the output of an algorithm. Technically, DP masks sensitive information by the addition of suitably scaled noise, thereby creating a trade-off between privacy (formalized as a privacy budget) and utility (measured by the expected accuracy of the estimates). This trade-off becomes particularly challenging when more than one quantity is meant to be estimated from the same private data, as is the case when estimating multiple data moments. Done naively, the privacy budget has to be split between the estimates, resulting in more noise and lower accuracy for both of them. In this work, we introduce a new method, Joint Moment Estimation (JME), that is able to privately estimate the first and second moments of vector-valued data without suffering from this shortcoming. Algorithmically, JME relies on the recent matrix factorization (MF) mechanism (Li et al., 2015) for continual release DP to individually compute the first and the second moments of the data, thereby making it flexible to accommodate a variety of settings. For example, besides the standard uniform sum or weighted average across data items, exponentially weighted averages or sliding-window estimates are readily possible. The key innovation of JME lies in our theoretical analysis of its properties. By jointly analyzing the sensitivity of the otherwise independent estimation processes, in combination with considering a carefully calibrated trade-off between them, we show that one can privately estimate the second-moment matrix without having to increase the amount of noise required to keep the first moment private. In this sense, we obtain privacy of the second moment for free. JME is practical and easily implemented using standard programming languages and toolboxes. We demonstrate this by showcasing two applications, one classical and one modern. First, we use JME for continual density estimation using a multivariate Gaussian model, e.g. we estimate the running mean and covariance matrix of a sequence of vector-valued observations. Our experiments confirm that JME achieves a lower Frobenius norm error for the covariance matrix in high-privacy regimes as well as a better fit to the true distribution as measured by the Kullback-Leibler divergence. Second, we integrate JME into the Adam optimizer, which is widely used in deep learning. Here, as well, we observe that JME achieves better optimization accuracy than baseline methods in high-privacy and small-batch-size regimes.
method symbol coefficients
prefix sum $E_{1}$ $\mathbbm{1}{\{i\leq t\}}$
exponential (weight $\beta$) $E_{\beta}$ $\beta^{t-i}\mathbbm{1}\{i\leq t\}$
standard average $V$ $\frac{1}{t}\mathbbm{1}\{i\leq t\}$
sliding window (size $k$) $W_{k}$ $\frac{1}{k}\mathbbm{1}\{t-k\,<\,i\,\leq\,t\}$
Table 1: Common workload matrices for moment estimation in (1).
@@ 2 Background
We study the problem of continually estimating the pair of weighted sums of the first and second moments in a differentially private manner. Namely, consider a sequence of $d$-dimensional vectors $x_{1},\dots,x_{n}\in\mathcal{X}$, where $\mathcal{X}=\{(x_{1},\dots,x_{n})\;:\;\max_{i}\|x_{i}\|_{2}\leq\zeta\wedge x_{i}\in\mathbb{R}^{d}\}$ for some fixed constant $\zeta>0$. At each step $t$, we aim to estimate the following pair of sums in a differentially private way:
@@
for arbitrary coefficients $a^{t}_{i}\in\mathbb{R}$ and $b^{t}_{i}\in\mathbb{R}$. This formulation includes many practical schemes, see Table 1.
@@
To express this problem compactly, we rewrite it in matrix form. We collect the coefficients into lower triangular workload matrices, $A_{1}=(a_{i}^{t})_{1\leq i,t\leq n}\in\mathbb{R}^{n\times n}$ and $A_{2}=(b_{i}^{t})_{1\leq i,t\leq n}\in\mathbb{R}^{n\times n}$. We stack the data as rows into a matrix $X\in\mathbb{R}^{n\times d}$. Then, all terms of the sums in (1) can then be expressed compactly [^1] For simplicity, we slightly abuse the notation and make $S$ a matrix of shape $n\times d^{2}$ instead of a tensor of size $n\times d\times d$. as $Y=A_{1}X$ and $S=A_{2}(X\operatorname{\!\bullet\!}X)$, where the second matrix consists of the data vectors stacked as rows, and $\bullet$ denotes the Face-Splitting Product or transposed Khatri-Rao product (Esteve et al., 2009), $(X\operatorname{\!\bullet\!}X)_{ij}:=(x_{i}\otimes x_{i})_{j}=\textsf{vec}(x_{i}x_{i}^{\top})_{j}$, where $\otimes$ is the Kronecker product of two vectors.
@@
To protect the privacy of individual data elements, we rely on the notion of differential privacy, considering neighboring datasets as those differing by a single data point; see e.g.Dwork & Roth (2014) for an introduction. Recently, a series of works have developed effective methods for privately estimating individual matrices in the aforementioned product form by means of the the matrix mechanism (Li et al., 2015; Choquette-Choo et al., 2024; Denisov et al., 2022; Choquette-Choo et al., 2023b; Dvijotham et al., 2024; Kalinin & Lampert, 2024; Henzinger et al., 2024; Henzinger & Upadhyay, 2025; Fichtenberger et al., 2023; Henzinger et al., 2023; Choquette-Choo et al., 2023a; Kairouz et al., 2021; McMahan et al., 2024). Its core insight is that, in the continual release setting, outputs at later steps should require less noise to be made private than earlier ones because more data points contributed to their computation. One can exploit this fact by adding noise that is correlated between the steps instead of being independent. This way, at each step, some of the noise added at an earlier stage can be removed again, thereby resulting in a less noisy result without lowering the privacy guarantees.
@@
Technically, the matrix mechanism relies on an invertible noise shaping matrix $C$. For our theoretical results, we often also assume that $C$ has decreasing column norms, As the private estimate of a product $AX$, one computes $\widehat{AX}=A(X+C^{-1}Z)$, where $Z$ is a matrix of i.i.d. Gaussian noise. This estimate is $(\epsilon,\delta)$-differentially private if the noise magnitude is at least $\text{sens}(CX)\cdot\sigma_{\epsilon,\delta}$, where $\sigma_{\epsilon,\delta}$ denotes the variance required in the Gaussian distribution to ensure $(\epsilon,\delta)$-DP for sensitivity $1$ queries [^2] Here and in the following we do not provide a formula for $\sigma_{\epsilon,\delta}$, because closed-form expressions have been shown to be suboptimal in some regimes. Instead, we recommend determining its value numerically, such as in (Balle & Wang, 2018). For reference, typical values of $\sigma_{\epsilon,\delta}$ lie between approximately $0.5$ (low privacy, e.g. $\epsilon=8,\delta=10^{-3})$ and $50$ (high privacy, e.g. $\epsilon=0.1,\delta=10^{-9})$.. $\text{sens}(\cdot)$ denotes the sensitivity, which, for any function $F$, is defined as
@@
where $X\sim X^{\prime}$ denote any two data matrices, that are neighboring, i.e. identical except for one of the data vectors.
@@ 3 Joint Moment Estimation (JME)
Algorithm 1 Joint Moment Estimation (JME)
0:  stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$
0:  workload matrices $A_{1},A_{2}\in\mathbb{R}^{n\times n}$
0:  privacy parameters $(\epsilon,\delta)$
0:  noise shaping matrices $C_{1},C_{2}\in\mathbb{R}^{n\times n}$ (invertible with decreasing column norm) default: $\text{I}_{n\times n},\text{I}_{n\times n}$
  $\lambda\leftarrow\|C_{1}\|_{1\to 2}^{2}/(c_{d}\zeta^{2}\|C_{2}\|_{1\to 2}^{2})$ // scaling parameter,   where $c_{1}=\frac{8}{11+5\sqrt{5}}$, and $c_{d}=2$ for $d\geq 2$
  $s\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // joint sensitivity
  $Z_{2}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d^{2}}$ // 2nd moment noise
  for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
  end for
@@
Note that privacy for free is a quite remarkable property of JME. Generally, when using the same data more than once for private computations, more noise for each of them is required to ensure the overall privacy, following, e.g., the composition theorems of DP (Kairouz et al., 2015). In some situations, however, one might also prefer a more flexible way to trade-off between the two moment estimates. For this, we present $\lambda$-JME in the appendix (Algorithm 3), a variant of JME that has $\lambda$ as a free hyper-parameter.
3.1 Properties of JME
For some inputs $X=(x_{1},\dots,x_{n})\in\mathbb{R}^{n\times d}$, let $Y=(Y_{1},\dots,Y_{n})\in\mathbb{R}^{n\times d}$ and $S=(S_{1},\dots,S_{n})\in\mathbb{R}^{n\times d^{2}}$ be the matrices of their first and second moments after each step, and let $\widehat{Y}=(Y_{1},\dots,Y_{n})\in\mathbb{R}^{n\times d}$ and $\widehat{S}=(S_{1},\dots,S_{n})\in\mathbb{R}^{n\times d^{2}}$ be the private estimates computed by Algorithm 1 with workload matrix $A$ and noise shaping matrix $C$. Then, the following properties hold.
@@ 3.1 Properties of JME
For any input $X$, the expected approximation error of Algorithm 1 for the first and second moments are
@@
$\sqrt{\mathbb{E}\|S-\widehat{S}\|^{2}_{\text{F}}}$ $=2\zeta\sqrt{c_{d}}d\sigma_{\epsilon,\delta}\|C_{2}\|_{1\to 2}\|A_{2}C_{2}^{-1}\|_{\text{F}},$ (4)
with $c_{d}$ as defined in Algorithm 1.
3.2 Proofs
@@
The statement follows from the fact that for any $t=1,\dots,n$, JME’s estimates $\widehat{x_{t}}$ and $\widehat{x_{t}\otimes x_{t}}$ of $x_{t}$ and $x_{t}\otimes x_{t}$ are unbiased with symmetric noise distribution because they are constructed by adding zero-mean Gaussian noise. ∎
Sketch of the Proof of Theorem 3.2.
By the properties of the Gaussian mechanism, it suffices to show that the overall sensitivity of jointly estimating both moments has at most the value $s=2\zeta\|C_{1}\|_{1\to 2}$, as stated in Algorithm 1.
@@
The following Lemma (proved in the appendix) characterizes the values of the joint sensitivity as a function of $\lambda$.
@@
Assume that the matrices $C_{1}$ and $C_{2}$ have norm-decreasing columns. Then, for any $\lambda>0$ holds:
@@
$r_{d}(\nu)=\begin{cases}\frac{1}{8}(3-\tau)^{2}(\nu\tau+1+\nu),&\text{if }\nu>\frac{11+5\sqrt{5}}{8},\,d=1,\\[5.0pt] 2+2\nu+\frac{1}{2\nu},&\text{if }\nu>\frac{1}{2},\,d>1,\\[5.0pt] 4,&\text{otherwise,}\end{cases}$ (7)
where $\tau=\sqrt{1-\frac{2}{\nu}}$.
Algorithm 1 uses the scaling parameter $\lambda=\|C_{1}\|_{1\to 2}^{2}/(c_{d}\zeta^{2}\|C_{2}\|_{1\to 2}^{2})$, so, by Lemma 3.5, the sensitivity of estimating both moment has the value $\sqrt{\zeta^{2}\|C_{1}\|_{1\to 2}^{2}r_{d}(c_{d})}$. The choice of $c_{d}$ implies that $r_{d}(c_{d})=4$, which establishes that $\text{sens}_{\lambda}(C_{1},C_{2})=s$, which concludes the proof of Theorem 3.2. ∎
Sketch of the Proof of Theorem 3.3.
@@
$\operatorname{\mathbb{E}}_{Z_{1}}\|\widehat{Y}-Y\|^{2}_{\text{F}}$ $=\|A_{1}C^{-1}_{1}\|^{2}_{\text{F}}\cdot\sigma^{2}_{\epsilon,\delta}s^{2}d$

### v2

@@ Abstract
We propose Joint Moment Estimation (JME), a method for continually and privately estimating both the first and second moments of a data stream with reduced noise compared to naive approaches. JME supports the matrix mechanism and exploits a joint sensitivity analysis to identify a privacy regime in which the second-moment estimation incurs no additional privacy cost, thereby improving accuracy while maintaining privacy. We demonstrate JME’s effectiveness in two applications: estimating the running mean and covariance matrix for Gaussian density estimation and model training with DP-Adam.
@@ 1 Introduction
Estimating the first and second moments of data is a fundamental step in many machine learning algorithms, ranging from foundational methods, such as linear regression, principal component analysis, or Gaussian model fitting, to state-of-the-art neural network components, such as BatchNorm (pmlr-v37-ioffe15, ), and optimizers, such as Adam (kingma2014adam, ). Often, these estimates must be computed and updated continuously, called the continual release setting (dwork2010differential, ). For example, this is the case when the data arrives sequentially and intermediate results are needed without delay, such as for sequential optimization algorithms like Adam, or for real-time systems in healthcare, finance, or recommendation systems. In such scenarios, ensuring the privacy of sensitive data, such as user information or medical records, is essential.
Differential Privacy (DP) is a widely established formal notion of privacy that ensures that the inclusion or exclusion of any single individual’s data has a limited impact on the output of an algorithm. Technically, DP masks sensitive information by adding suitably scaled noise, thereby creating a trade-off between privacy (formalized as a privacy budget) and utility (measured by the expected accuracy of the estimates). This trade-off becomes particularly challenging when more than one quantity is meant to be estimated from the same private data, as is the case when estimating multiple data moments. Done naively, the privacy budget has to be split between the estimates, resulting in more noise and lower accuracy for both of them.
@@
In this work, we introduce a new method, Joint Moment Estimation (JME), that is able to privately estimate the first and second moments of vector-valued data without suffering from this shortcoming. Algorithmically, JME relies on the recent matrix factorization (MF) mechanism (li2015matrix, ) for continual release DP to individually compute the first and the second moments of the data, thereby making it flexible to accommodate a variety of settings. For example, besides the standard uniform sum or weighted average across data items, exponentially weighted averages or sliding-window estimates are readily possible. The key innovation of JME lies in our theoretical analysis of its properties. By jointly analyzing the sensitivity of the otherwise independent estimation processes, in combination with considering a carefully calibrated trade-off between them, we show that one can estimate privately the second-moment matrix without having to increase the amount of noise required to keep the first moment private. In this sense, we obtain privacy of the second moment for free.
JME is practical and easily implemented using standard programming languages and toolboxes. We demonstrate this by showcasing two applications, one classical and one modern. First, we use JME for continual density estimation using a multivariate Gaussian model, e.g. we estimate the running mean and covariance matrix of a sequence of vector-valued observations. Our experiments confirm that JME achieves a lower Frobenius norm error for the covariance matrix in high-privacy regimes as well as a better fit to the true distribution as measured by the Kullback-Leibler divergence. Second, we integrate JME into the Adam optimizer, which is widely used in deep learning. Here, as well, we observe that JME achieves better optimization accuracy than baseline methods in high-privacy and small-batch-size regimes.
@@ 2 Background
We study the problem of differentially private continually estimation of the pair of weighted sums of the first and second moments, see Section 6 for a discussion of related works. Specifically, consider a sequence of $d$-dimensional vectors $x_{1},\dots,x_{n}\in\mathcal{X}$, where $\mathcal{X}=\{(x_{1},\dots,x_{n})\;:\;x_{i}\in\mathbb{R}^{d}\ \wedge\ \max_{i}\|x_{i}\|_{2}\leq\zeta\}$ for some fixed constant $\zeta>0$. At each step $t$, we aim to privately estimate the following pair of sums:
@@
for arbitrary coefficients $a^{t}_{i}\in\mathbb{R}$ and $b^{t}_{i}\in\mathbb{R}$. This formulation includes many practical schemes; see Table 1 in the appendix. To express this problem compactly, we rewrite it in matrix form. We collect the coefficients into lower triangular workload matrices, $A_{1}=(a_{i}^{t})_{1\leq i,t\leq n}\in\mathbb{R}^{n\times n}$ and $A_{2}=(b_{i}^{t})_{1\leq i,t\leq n}\in\mathbb{R}^{n\times n}$. We stack the data as rows into a matrix $X\in\mathbb{R}^{n\times d}$. Then, all terms of the sums in (1) can then be expressed compactly [^1] We slightly abuse the notation and make $S$ a matrix of shape $n\times d^{2}$ instead of a tensor of size $n\times d\times d$. as $Y=A_{1}X$ and $S=A_{2}(X\operatorname{\bullet}X)$, where the second matrix consists of the data vectors stacked as rows, and $\bullet$ denotes the Face-Splitting Product or transposed Khatri-Rao product (esteve2009, ), $(X\operatorname{\bullet}X)_{ij}:=(x_{i}\otimes x_{i})_{j}=\textsf{vec}(x_{i}x_{i}^{\top})_{j}$, where $\otimes$ is the Kronecker product of two vectors.
@@
To protect the privacy of individual data elements, we rely on the notion of differential privacy, considering neighboring datasets as those differing by a single data point; see e.g.(dwork2014algorithmic, ) for an introduction. Recently, a series of works have developed effective methods for privately estimating individual matrices in the aforementioned product form by means of the matrix mechanism (li2015matrix, ; CNPBINDPL, ; Denisov, ; MultiEpoch, ; dvijotham2024efficient, ; kalinin2024, ; henzinger2024, ). Its core insight is that, in the continual release setting, outputs at later steps should require less noise to be made private than earlier ones because more data points contributed to their computation. One can exploit this fact by adding noise that is correlated between the steps instead of being independent. This way, at each step, some of the noise added at an earlier stage can be removed again, thereby resulting in a less noisy result without lowering the privacy guarantees.
@@
Technically, the matrix mechanism relies on an invertible noise shaping matrix $C$. For our theoretical results, we often also assume that $C$ has decreasing column norms, As the private estimate of a product $AX$, one computes $\widehat{AX}=A(X+C^{-1}Z)$, where $Z$ is a matrix of i.i.d. Gaussian noise. This estimate is $(\epsilon,\delta)$-differentially private if the noise magnitude is at least $\text{sens}(CX)\cdot\sigma_{\epsilon,\delta}$, where $\sigma_{\epsilon,\delta}$ denotes the variance required in the Gaussian distribution to ensure $(\epsilon,\delta)$-DP for sensitivity $1$ queries [^2] Here and in the following we do not provide a formula for $\sigma_{\epsilon,\delta}$, because closed-form expressions have been shown to be suboptimal in some regimes. Instead, we recommend determining its value numerically, such as in (balle2018improving, ). For reference, typical values of $\sigma_{\epsilon,\delta}$ lie between approximately $0.5$ (low privacy, e.g. $\epsilon=8,\delta=10^{-3})$ and $50$ (high privacy, e.g. $\epsilon=0.1,\delta=10^{-9})$.. $\text{sens}(\cdot)$ denotes the sensitivity, which, for any function $F$, is defined as
@@
where $X\sim X^{\prime}$ denote two neighboring data matrices, i.e. identical except for one of the data vectors.
@@ 3 Joint Moment Estimation (JME)
Algorithm 1 Joint Moment Estimation (JME)
stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$,
workload matrices $A_{1},A_{2}\in\mathbb{R}^{n\times n}$
privacy parameters $(\epsilon,\delta)$
noise shaping matrices $C_{1},C_{2}\in\mathbb{R}^{n\times n}$ (invertible, decreasing column norm), defaults: Id
$\lambda\leftarrow\|C_{1}\|_{1\to 2}^{2}/(c_{d}\zeta^{2}\|C_{2}\|_{1\to 2}^{2})$ with $c_{1}=\frac{8}{11+5\sqrt{5}}$, and $c_{d}=2$ for $d\geq 2$  // scaling parameter
$s\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // joint sensitivity
$Z_{1}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d}$, $Z_{2}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d^{2}}$ // 1st and 2nd moment noise
for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$ and $\widehat{x_{t}\otimes x_{t}}\leftarrow x_{t}\otimes x_{t}+\lambda^{-1/2}[C_{2}^{-1}Z_{2}]_{[t,\cdot,\cdot]}$
end for
Note that privacy for free is a quite remarkable property of JME. Generally, when using the same data more than once for private computations, more noise for each of them is required to ensure the overall privacy, following, e.g., the composition theorems of DP (kairouz15, ). In some situations, however, one might also prefer a more flexible way to trade-off between the two moment estimates. For this, we present $\lambda$-JME in the appendix (Algorithm 3), a variant of JME that has $\lambda$ as a free hyper-parameter.
Properties of JME
For some inputs $X=(x_{1},\dots,x_{n})\in\mathbb{R}^{n\times d}$, let $Y=(Y_{1},\dots,Y_{n})\in\mathbb{R}^{n\times d}$ and $S=(S_{1},\dots,S_{n})\in\mathbb{R}^{n\times d^{2}}$ be the matrices of their first and second moments after each step, and let $\widehat{Y}=(Y_{1},\dots,Y_{n})\in\mathbb{R}^{n\times d}$ and $\widehat{S}=(S_{1},\dots,S_{n})\in\mathbb{R}^{n\times d^{2}}$ be the private estimates computed by Algorithm 1 with workload matrix $A$ and noise shaping matrix $C$. Then, we have the following.
@@
For any input $X$, let $c_{d}$ be as defined in Algorithm 1. Then the expected approximation error of Algorithm 1 for the first and second moments are
@@
$\sqrt{\mathbb{E}\|S-\widehat{S}\|^{2}_{\text{F}}}$ $=2\zeta^{2}\sqrt{c_{d}}d\sigma_{\epsilon,\delta}\|C_{2}\|_{1\to 2}\|A_{2}C_{2}^{-1}\|_{\text{F}}.$ (4)
@@
The statement follows from the fact that for any $t=1,\dots,n$, JME’s estimates $\widehat{x_{t}}$ of $x_{t}$ and $\widehat{x_{t}\otimes x_{t}}$ of $x_{t}\otimes x_{t}$ are unbiased with symmetric noise distribution because they are constructed by adding zero-mean Gaussian noise. ∎
Proof sketch of Theorem 3.2.
The result follows using Gaussian mechanism once we show that the overall sensitivity of jointly estimating both moments has at most $s=2\zeta\|C_{1}\|_{1\to 2}$, as stated in Algorithm 1.
@@
The following lemma (shown in the appendix) characterizes the values of the joint sensitivity as a function of $\lambda$.
@@
Assume that the matrices $C_{1}$ and $C_{2}$ have norm-decreasing columns [^3] The decreasing column norm structure ensures that the optimum is simultaneously achieved for the first columns of the matrices $C_{1}$ and $C_{2}$. This can be relaxed by solving a maximization problem over the column indices, which could be done numerically. Moreover, the matrices that we have in mind for practical tasks are Toeplitz, for which the norm-decreasing property is fulfilled naturally.. Then, for any $\lambda>0$ holds:
@@
$r_{d}(\nu)=\begin{cases}\frac{1}{8}(3-\tau)^{2}(\nu\tau+1+\nu),\text{ with }\tau=\sqrt{1-\frac{2}{\nu}},\quad&\text{if }\nu>\frac{11+5\sqrt{5}}{8},\,d=1,\\ 2+2\nu+\frac{1}{2\nu},&\text{if }\nu>\frac{1}{2},\,d>1,\\[5.0pt] 4,&\text{otherwise.}\end{cases}$ (7)
Algorithm 1 uses the scaling parameter $\lambda=\|C_{1}\|_{1\to 2}^{2}/(c_{d}\zeta^{2}\|C_{2}\|_{1\to 2}^{2})$, so, by Lemma 3.5, the sensitivity of estimating both moment has the value $\sqrt{\zeta^{2}\|C_{1}\|_{1\to 2}^{2}r_{d}(c_{d})}$. The choice of $c_{d}$ implies that $r_{d}(c_{d})=4$, which yields $\text{sens}_{\lambda}(C_{1},C_{2})=s$. This concludes the proof of Theorem 3.2. ∎
Proof sketch of Theorem 3.3.
@@
$\operatorname{\mathbb{E}}_{Z_{1}}\|\widehat{Y}-Y\|^{2}_{\text{F}}$ $=\|A_{1}C^{-1}_{1}\|^{2}_{\text{F}}\cdot\sigma^{2}_{\epsilon,\delta}s^{2}d$ (8)